プリンキピア

第5章連続確率変数

5.1 Continuous Probability Functions5.1 連続確率関数

We begin by defining a continuous probability density function. We use the function notation f(x). Intermediate algebra may have been your first formal introduction to functions. In the study of probability, the functions we study are special. We define the function f(x) so that the area between it and the x-axis is equal to a probability. Since the maximum probability is one, the maximum area is also one. For continuous probability distributions, PROBABILITY = AREA.
英語のヒント
まず連続確率密度関数を定義する。関数の表記f(x)を用いる。関数を初めて正式に学んだのは、代数の中級課程だったかもしれない。確率の学習で扱う関数は特別なものである。関数f(x)とx軸との間の面積が確率に等しくなるように、f(x)を定義する。確率の最大値は1なので、面積の最大値も1である。連続確率分布では、確率 = 面積である。
Example 5.1例題 5.1
Consider the function f(x) = 120 for 0 ≤ x ≤ 20. x = a real number. The graph of f(x) = 120 is a horizontal line. However, since 0 ≤ x ≤ 20, f(x) is restricted to the portion between x = 0 and x = 20, inclusive.
英語のヒント
0 ≤ x ≤ 20で関数f(x) = 120を考える。xは実数である。f(x) = 120のグラフは水平な直線である。ただし、0 ≤ x ≤ 20なので、f(x)はx = 0からx = 20までの両端を含む部分に限定される。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸びる。x軸から端点(20, 1/20)へ縦線を引き、長方形を作っている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.5図 5.5
f(x) = 120 for 0 ≤ x ≤ 20.
0 ≤ x ≤ 20で、f(x) = 120。
The graph of f(x) = 120 is a horizontal line segment when 0 ≤ x ≤ 20.
0 ≤ x ≤ 20のとき、f(x) = 120のグラフは水平な線分である。
The area between f(x) = 120 where 0 ≤ x ≤ 20 and the x-axis is the area of a rectangle with base = 20 and height = 120.
0 ≤ x ≤ 20におけるf(x) = 120とx軸との間の面積は、底辺20、高さ120の長方形の面積である。
AREA=20( 1 20 )=1
面積=20( 1 20 )=1
Suppose we want to find the area between f(x) = 120 and the x-axis where 0 < x < 2.
0 < x < 2の範囲で、f(x) = 120とx軸との間の面積を求めたいとする。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸びる。x軸から端点(20, 1/20)へ縦線を引いて長方形を作り、その中のx = 0からx = 2までを着色している。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.6図 5.6
AREA = (2 – 0)( 1 20 ) = 0.1
面積 = (2 – 0)( 1 20 ) = 0.1
(2 – 0) = 2 = base of a rectangle
(2 – 0) = 2 = 長方形の底辺
The area corresponds to a probability. The probability that x is between zero and two is 0.1, which can be written mathematically as P(0 < x < 2) = P(x < 2) = 0.1.
この面積は確率に対応する。xが0と2の間にある確率は0.1であり、数学的にはP(0 < x < 2) = P(x < 2) = 0.1と書ける。
Suppose we want to find the area between f(x) = 120 and the x-axis where 4 < x < 15.
4 < x < 15の範囲で、f(x) = 120とx軸との間の面積を求めたいとする。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸びる。x軸から端点(20, 1/20)へ縦線を引いて長方形を作り、その中のx = 4からx = 15までを着色している。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.7図 5.7
AREA = (15 – 4)( 1 20 ) = 0.55
面積 = (15 – 4)( 1 20 ) = 0.55
(15 – 4) = 11 = the base of a rectangle
(15 – 4) = 11 = 長方形の底辺
The area corresponds to the probability P(4 < x < 15) = 0.55.
この面積は、確率P(4 < x < 15) = 0.55に対応する。
Suppose we want to find P(x = 15). On an x-y graph, x = 15 is a vertical line. A vertical line has no width (or zero width). Therefore, P(x = 15) = (base)(height) = (0) ( 1 20 ) = 0
P(x = 15)を求めたいとする。xy平面では、x = 15は垂直な直線である。垂直な直線には幅がない(幅が0である)。したがって、P(x = 15) = (底辺)(高さ)= (0) ( 1 20 ) = 0となる。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸び、x軸から端点(20, 1/20)への縦線とともに長方形を作る。x = 15でも横軸からグラフへ縦線が伸びている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.8図 5.8
P(X ≤ x), which can also be written as P(X < x) for continuous distributions, is called the cumulative distribution function or CDF. Notice the "less than or equal to" symbol. We can also use the CDF to calculate P(X > x). The CDF gives "area to the left" and P(X > x) gives "area to the right." We calculate P(X > x) for continuous distributions as follows: P(X > x) = 1 – P (X < x).
P(X ≤ x)を累積分布関数、またはCDFと呼ぶ。連続分布ではP(X < x)とも書ける。「以下」の記号に注意しよう。CDFを使ってP(X > x)を計算することもできる。CDFは「左側の面積」を、P(X > x)は「右側の面積」を与える。連続分布では、P(X > x) = 1 − P(X < x)として計算する。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸び、x軸から端点(20, 1/20)への縦線とともに長方形を作る。ある値xより左の領域が着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.9図 5.9
Label the graph with f(x) and x. Scale the x and y axes with the maximum x and y values. f(x) = 1 20 , 0 ≤ x ≤ 20.
グラフにf(x)とxを記入せよ。xとyの最大値を用いて、両軸に目盛りを付けよ。f(x) = 1 20 、0 ≤ x ≤ 20。
To calculate the probability that x is between two values, look at the following graph. Shade the region between x = 2.3 and x = 12.7. Then calculate the shaded area of a rectangle.
xが二つの値の間にある確率を計算するため、次のグラフを見よう。x = 2.3とx = 12.7の間を着色し、長方形の着色部分の面積を求めよ。
関数f(x) = 1/20のグラフ。点(0, 1/20)から点(20, 1/20)まで水平線が伸び、x軸から端点(20, 1/20)への縦線とともに長方形を作る。長方形内のx = 2.3からx = 12.7までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.10図 5.10
P(2.3<x<12.7)=(base)(height)=(12.7−2.3)( 1 20 )=0.52
P(2.3<x<12.7)=(底辺)(高さ)=(12.7−2.3)( 1 20 )=0.52

Chapter Review章の復習

The probability density function (pdf) is used to describe probabilities for continuous random variables. The area under the density curve between two points corresponds to the probability that the variable falls between those two values. In other words, the area under the density curve between points a and b is equal to P(a < x < b). The cumulative distribution function (cdf) gives the probability as an area. If X is a continuous random variable, the probability density function (pdf), f(x), is used to draw the graph of the probability distribution. The total area under the graph of f(x) is one. The area under the graph of f(x) and between values a and b gives the probability P(a < x < b).
確率密度関数(pdf)は、連続確率変数の確率を記述するために用いる。二点の間の密度曲線の下の面積は、変数がその二つの値の間に入る確率に対応する。つまり、点aとbの間の密度曲線の下の面積はP(a < x < b)に等しい。累積分布関数(cdf)は、確率を面積として与える。Xが連続確率変数なら、確率密度関数f(x)で確率分布のグラフを描く。f(x)のグラフの下の総面積は1である。値aとbの間にあるf(x)のグラフの下の面積が、確率P(a < x < b)を与える。
左のグラフは一般的な密度曲線y = f(x)を示す。曲線より下でx軸より上の領域が着色され、面積は1である。これはすべての可能な結果が曲線で表されることを示す。右も同じ密度曲線で、x = aとx = bの縦線が軸から曲線まで伸び、その間が着色されている。着色部分の面積は、値xがaとbの間にある確率を表す。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.11図 5.11
The cumulative distribution function (cdf) of X is defined by P (X ≤ x). It is a function of x that gives the probability that the random variable is less than or equal to x.
Xの累積分布関数(cdf)はP(X ≤ x)で定義される。これはxの関数で、確率変数がx以下となる確率を与える。

Formula Review公式の復習

Probability density function (pdf) f(x):
確率密度関数(pdf)f(x):
  • f(x) ≥ 0
  • The total area under the curve f(x) is one.
    曲線f(x)の下の総面積は1である。
Cumulative distribution function (cdf): P(X ≤ x)
累積分布関数(cdf):P(X ≤ x)
Exercises 1–15練習問題 1–15
Exercise 1問題 1
Which type of distribution does the graph illustrate?
このグラフはどの種類の分布を表しているか。
横軸は0から10。分布はx = 3からx = 8までの長方形で表される。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.12図 5.12
Solution解き方
Uniform Distribution
一様分布
Exercise 2問題 2
Which type of distribution does the graph illustrate?
このグラフはどの種類の分布を表しているか。
右下がりのグラフ。y軸上の点から始まり、右端に向かってx軸に近づく。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.13図 5.13
Exercise 3問題 3
Which type of distribution does the graph illustrate?
このグラフはどの種類の分布を表しているか。
釣鐘型のグラフ。左右対称で、x = 0で最も高くなり、頂点の両側で緩やかに下がってx軸に近づく。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.14図 5.14
Solution解き方
Normal Distribution
正規分布
Exercise 4問題 4
What does the shaded area represent? P(___< x < ___)
着色部分の面積は何を表すか。P(___ < x < ___)
一様分布のグラフ。横軸は0から10で、分布はx = 1からx = 8までの長方形で表される。長方形内のx = 2からx = 5までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.15図 5.15
Exercise 5問題 5
What does the shaded area represent? P(___< x < ___)
着色部分の面積は何を表すか。P(___ < x < ___)
指数分布のグラフ。y軸上の点から右下がりに伸び、右端に向かってx軸に近づく。グラフの下のx = 6からx = 7までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.16図 5.16
Solution解き方
P(6 < x < 7)
Exercise 6問題 6
For a continuous probablity distribution, 0 ≤ x ≤ 15. What is P(x > 15)?
ある連続確率分布で、0 ≤ x ≤ 15である。P(x > 15)はいくらか。
Exercise 7問題 7
What is the area under f(x) if the function is a continuous probability density function?
関数が連続確率密度関数であるとき、f(x)の下の面積はいくらか。
Solution解き方
one
1
Exercise 8問題 8
For a continuous probability distribution, 0 ≤ x ≤ 10. What is P(x = 7)?
ある連続確率分布で、0 ≤ x ≤ 10である。P(x = 7)はいくらか。
Exercise 9問題 9
A continuous probability function is restricted to the portion between x = 0 and 7. What is P(x = 10)?
ある連続確率関数は、x = 0から7までの部分に限定されている。P(x = 10)はいくらか。
Solution解き方
zero
0
Exercise 10問題 10
f(x) for a continuous probability function is 1 5 , and the function is restricted to 0 ≤ x ≤ 5. What is P(x < 0)?
連続確率関数f(x)は 1 5 で、0 ≤ x ≤ 5に限定されている。P(x < 0)はいくらか。
Exercise 11問題 11
f(x), a continuous probability function, is equal to 1 12 , and the function is restricted to 0 ≤ x ≤ 12. What is P (0 < x < 12)?
連続確率関数f(x)は 1 12 に等しく、0 ≤ x ≤ 12に限定されている。P(0 < x < 12)はいくらか。
Solution解き方
one
1
Exercise 12問題 12
Find the probability that x falls in the shaded area.
Xが陰影を付けた領域に入る確率を求めよ。
一様分布のpdfである関数f(x) = 1/9のグラフ。点(0, 1/9)から点(9, 1/9)まで水平線が伸びる。x = 9でx軸からグラフへ縦線を引き、二つの座標軸を辺とする長方形を作る。x = 6からx = 8までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.17図 5.17
Exercise 13問題 13
Find the probability that x falls in the shaded area.
Xが陰影を付けた領域に入る確率を求めよ。
一様分布のpdfである関数f(x) = 1/8のグラフ。点(0, 1/8)から点(8, 1/8)まで水平線が伸びる。x = 8でx軸からグラフへ縦線を引き、二つの座標軸を辺とする長方形を作る。x = 0からx = 5までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.18図 5.18
Solution解き方
0.625
Exercise 14問題 14
Find the probability that x falls in the shaded area.
Xが陰影を付けた領域に入る確率を求めよ。
一様分布のpdfである関数f(x) = 1/10のグラフ。点(0, 1/10)から点(10, 1/10)まで水平線が伸びる。x = 10でx軸からグラフへ縦線を引き、二つの座標軸を辺とする長方形を作る。x = 2.5からx = 5.5までが着色されている。

Copyright Rice University, OpenStax, Introductory Statistics 2e, CC BY 4.0. · CC BY 4.0 · Source出典

Figure 5.19図 5.19
Exercise 15問題 15
f(x), a continuous probability function, is equal to 1 3 and the function is restricted to 1 ≤ x ≤ 4. Describe P( x> 3 2 ).
連続確率関数f(x)は 1 3 に等しく、1 ≤ x ≤ 4に限定されている。 P( x> 3 2 ) を説明せよ。
Solution解き方
The probability is equal to the area from x = 3 2 to x = 4 above the x-axis and up to f(x) = 1 3 .
確率は、x = 3 2 からx = 4までの、x軸より上でf(x) = 1 3 までの面積に等しい。